OpenStudy (anonymous):

Rationalize the denominator of square root 50 / square root 5

jimthompson5910 (jim_thompson5910):

Hint: Use the idea that \[\Large \frac{\sqrt{x}}{\sqrt{y}}=\sqrt{\frac{x}{y}}\]

OpenStudy (anonymous):

5 square root of 5 / 2 ?

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

So use the idea I gave you above to go from \[\Large \frac{\sqrt{50}}{\sqrt{5}}\] to \[\Large \sqrt{\frac{50}{5}}\]

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

ohhh square root 10

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

nice job

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome